# Hadronic scattering in (1+1)D SU(2) lattice gauge theory from tensor networks

### v2

@@ 1 Introduction
In the context of particle scattering, recent efforts have achieved the first simulations of elastic and inelastic processes in Abelian gauge theories using tensor-network and quantum-simulation techniques, see e.g. Papaefstathiou:2024zsu ; Chai:2023qpq ; Bennewitz:2025nhz ; Davoudi:2025rdv ; Farrell:2025nkx ; Barata:2025hgx ; Abel:2025zxb ; Carena:2024peb ; Jha:2024jan ; Zemlevskiy:2024vxt ; Ingoldby:2025bdb ; Schuhmacher:2025ehh ; Su:2024uuc , and also Pavesic:2025nwm ; Cobos:2025krn ; DiMarcantonio:2025cmf ; Gonzalez-Cuadra:2024xul ; Joshi:2025rha ; Xu:2025abo for related extensions to (2+1)D models. However, so far the focus has been mainly restricted to Abelian theories. Extending this program to non-Abelian models represents a crucial step toward capturing phenomena such as confinement and baryon formation, that are central to QCD-like dynamics.
@@ 2 SU(2) lattice gauge theory in (1+1)D
$+\frac{ag^{2}}{2}\frac{3}{8}\sum_{n=1}^{N-1}(N-n)\left[1-2\sigma^{z}_{2n-1}-2\sigma^{z}_{2n}\right]$
@@
$+\frac{1}{16}\sum_{n=1}^{N-2}\sum_{m=n+1}^{N-1}(N-m)(\sigma^{z}_{2n-1}-\sigma^{z}_{2n})(\sigma^{z}_{2m-1}-\sigma^{z}_{2m})\,.$ (8)
@@
Figure 2: Meson to baryon mass ratio at evaluated with $N=60$ qubits compared with the analytical strong coupling estimation. Here $x=1/(ga)^{2}$ and the strong coupling limit corresponds to $1/x\gg 1$.
@@ 4 Numerical results
In this section, we present the numerical results for the simulation of scattering in the SU(2) lattice model using MPS. Following the above discussion, we divide our results into different baryon number sectors, $B=0,1,2$. For all the simulations, we use the same width $\sigma=4$ for the initial wavepackets.
@@ 4.2 $B=1$, meson-baryon scattering
From Fig. 7 one can observe that in both scenarios the meson state is delocalized, with part of it being reflected and the remaining tunnelling over the baryon state. This is clearly seen when computing the meson and baryon projections, which show the baryon to remain mainly unchanged, while the meson probability is spread in space. More, considering the internal distribution of the baryon number inside these states, we observe that the characteristic pattern of the two sates interfere constructively. Nonetheless, the entropy evolution is rather distinct from the $B=0$ sector, where in the unequal momentum scenario, after the scattering, the entropy peak survives around the scattering point. This indicates that, for the simulated times, the two wavefunctions overlap significantly and form a single state. Finally, in Fig. 8 we show the corresponding results using the information lattice. This detailed mapping of the correlations in the system supports the previous statements: the baryon state remains localized along its classical trajectory, while the meson completely overlaps with the latter. Notice that now new information levels $\ell$ are populated (at least not significantly), indicating that there is no formation of a new state with a different correlation profile, see discussion in Artiaco:2025qqq .

### v3

@@ 1 Introduction
In the context of particle scattering, recent efforts have achieved the first simulations of elastic and inelastic processes in Abelian gauge theories using tensor-network and quantum-simulation techniques, see e.g. Papaefstathiou:2024zsu ; Chai:2023qpq ; Bennewitz:2025nhz ; Davoudi:2025rdv ; Farrell:2025nkx ; Barata:2025hgx ; Abel:2025zxb ; Carena:2024peb ; Jha:2024jan ; Zemlevskiy:2024vxt ; Ingoldby:2025bdb ; Schuhmacher:2025ehh ; Su:2024uuc ; Belyansky:2023rgh , and also Pavesic:2025nwm ; Cobos:2025krn ; DiMarcantonio:2025cmf ; Gonzalez-Cuadra:2024xul ; Joshi:2025rha ; Xu:2025abo for related extensions to (2+1)D models. However, so far the focus has been mainly restricted to Abelian theories. Extending this program to non-Abelian models represents a crucial step toward capturing phenomena such as confinement and baryon formation, that are central to QCD-like dynamics.
@@ 2 SU(2) lattice gauge theory in (1+1)D
$+\frac{ag^{2}}{2}\Bigg\{\frac{3}{8}\sum_{n=1}^{N-1}(N-n)\left[1-2\sigma^{z}_{2n-1}-2\sigma^{z}_{2n}\right]$
@@
$+\frac{1}{16}\sum_{n=1}^{N-2}\sum_{m=n+1}^{N-1}(N-m)(\sigma^{z}_{2n-1}-\sigma^{z}_{2n})(\sigma^{z}_{2m-1}-\sigma^{z}_{2m})\Bigg\}\,.$ (8)
@@
Figure 2: Meson to baryon mass ratio at evaluated with $N=60$ qubits compared with the analytical strong coupling estimation. Here $x=1/(ga)^{2}$ and the strong coupling limit corresponds to $1/x\gg 1$. Uncertainty bar is computed from the absolute different between simulation results using $\chi_{\mathrm{max}}=80$ and $\chi_{\mathrm{max}}=200$.
@@ 3 Scattering simulation using tensor networks
The wavepacket operators are explicitly defined, up to a normalizing factor, as
@@ 4 Numerical results
In this section, we present the numerical results for the simulation of scattering in the SU(2) lattice model using MPS. Following the above discussion, we divide our results into different baryon number sectors, $B=0,1,2$. The simulation parameters for both the state preparation and real-time simulation are explicitly given in Section 3.
@@ 4.2 $B=1$, meson-baryon scattering
From Fig. 7 one can observe that in both scenarios the meson state is delocalized, with part of it being reflected and the remaining tunnelling over the baryon state. This is clearly seen when computing the meson and baryon projections, which show the baryon to remain mainly unchanged, while the meson probability is spread in space. More, considering the internal distribution of the baryon number inside these states, we observe that the characteristic pattern of the two sates interfere constructively. Nonetheless, the entropy evolution is rather distinct from the $B=0$ sector, where in the unequal momentum scenario, after the scattering, the entropy peak survives around the scattering point. This indicates that, for the simulated times, the two wavefunctions overlap significantly and form a single state. Note that the difference between $B=0$ and $B=1$ case is mostly connected to the fact that the scattering states are not symmetric rather than the non-Abelian nature of the of the setup. Finally, in Fig. 8 we show the corresponding results using the information lattice. This detailed mapping of the correlations in the system supports the previous statements: the baryon state remains localized along its classical trajectory, while the meson completely overlaps with the latter. Notice that now new information levels $\ell$ are populated (at least not significantly), indicating that there is no formation of a new state with a different correlation profile, see discussion in Artiaco:2025qqq .
@@ 4.3 $B=2$, baryon-baryon scattering
4.4 Numerical convergence
In this subsection, we provide a detailed demonstration for the numerical convergence in our real-time scattering simulation result using tensor network. As we already demonstrated convergence in the meson and hadron masses (see Fig. 2), we now focus on the numerical benchmark in the real-time simulation using TDVP algorithm. In particular, we use the two-site entanglement entropy $S_{n}(t)$ as the main target observable to quantify the error and convergence associated with the simulation.
In Fig. 10, we show the entanglement entropy $S_{n}(t)$ for different maximum bond dimensions $\chi_{\mathrm{max}}$ in the $B=0$, $B=1$, and $B=2$ scattering. We can see that $S_{n}(t)$ at selected time instances before, during, and after the scattering shows no qualitative dependences on the maximum bond dimensions. The associated error $|\Delta S_{n}(t)|=|S_{n}(t;\chi_{\mathrm{max}}=80)-S_{n}(t;\chi_{\mathrm{max}}=60)|$ calculated from $S_{n}(t)$ at two different $\chi_{\mathrm{max}}$ in the TDVP simulation is small, around $10^{-3}$, which is less than around 0.1% compared to the entropy itself, though slightly increased error is observed for a longer simulation time. In summary, our choice of maximum bond dimension $\chi_{\mathrm{max}}=80$ used in the main result of the manuscript is sufficiently large enough and the same convergence pattern has been observed for other observables as well.
Figure 10: Two-site entanglement entropy $S_{n}(t)$ for $B=0$, $B=1$, and $B=2$ versus two different maximum bond dimension $\chi_{\mathrm{max}}=60$ and 8080 used in TDVP. In the right column, $|\Delta S_{n}(t)|=|S_{n}(t;\chi_{\mathrm{max}}=80)-S_{n}(t;\chi_{\mathrm{max}}=60)|$.
Next, we investigate our choice of time step used in the real-time evolution. In Fig. 11, we show the entanglement entropy $S_{n}(t)$ for different time step $\alpha\delta t$ in the $B=0$, $B=1$, and $B=2$ scattering. Similarly, we see that $S_{n}(t)$ at selected time instances before, during, and after the scattering are not significantly affected by our choice of the time step, suggesting that $a\delta t=0.1$ is sufficiently small in our setup. We define the associated error $|\delta S_{n}(t)|=|S_{n}(t;a\delta t=0.1)-S_{n}(t;a\delta t=0.2)|$ to quantify the error from time step and $|\delta S_{n}(t)|\lesssim 10^{-3}$ is observed throughout the evolution.
Figure 11: Two-site entanglement entropy $S_{n}(t)$ for $B=0$, $B=1$, and $B=2$ versus two different time step $a\delta t=0.1$ and $0.2$ used in TDVP. In the right column, $|\delta S_{n}(t)|=|S_{n}(t;a\delta t=0.1)-S_{n}(t;a\delta t=0.2)|$.
Last but not the least, it is also interesting and important to investigate the finite volume effects for the numerical simulation. The volume or system size of the setup is $\mathcal{V}=Nag$, where $N$ is the number of qubits, $a$ the lattice spacing, and $g$ the coupling strength. Currently our volume is fixed at $\mathcal{V}=300$ in the state preparation and in the real-time simulation. To truly investigate the finite volume effect on the real-time simulation as well as on the mass spectrum, one needs to increase $N$ and extrapolate the numerical results as $x\rightarrow\infty$, whose procedure is well demonstrated in Ref. Banuls:2013jaa for the U(1) Schwinger model. Similar, one also need to follow the same procedure to truly investigate finite volume effect for the real-time simulation. However, the SU(2) Hamiltonian employed in this work is much more difficult due to its non-Abelian nature, resulting in having more complicated internal structure and higher-order interactions. Rigorous study on the finite size effect goes beyond the scope of this work and probably is best suited to a different representation with gauge field truncation, and therefore we leave it for for the future.